OpenStudy (anonymous):

Please Help! The rational function has a y-intercept. What is the equation for this function?

OpenStudy (anonymous):

Heres the graph! I can't figure out how to get the equation! I know how to do the reverse, use the function to get the graph! Please help!

OpenStudy (anonymous):

*The rational function has a y-intercept of 7, I meant to say at the top

OpenStudy (zzr0ck3r):

what does this function "look" like to you?

OpenStudy (zzr0ck3r):

you should know what the basic functions look like

OpenStudy (anonymous):

Anybody have any clue how to get the equation through the graph?? My textbook only shows how to use a function to find the graph... I mean what steps do i take to do it, i know what a function looks like

OpenStudy (zzr0ck3r):

you need to identify what parent function we are working with

OpenStudy (zzr0ck3r):

does this look like y=x, y=x^2,y=x^3, y = 1/x, .....?

OpenStudy (anonymous):

using general form y= (a/x-h) +k then translate?? like horizontal and vertical, the book refers to graphing rational functions like y=x^2+x-12/x^2-4. And finding graph but how do i find the equation?

OpenStudy (zzr0ck3r):

(a/x-h) +k

OpenStudy (zzr0ck3r):

ok so, the h is the horizontal shift of 1/x

OpenStudy (zzr0ck3r):

which way is your graph shifted?

OpenStudy (anonymous):

2 units to the left

OpenStudy (zzr0ck3r):

ok so we want (a/x+2) +k

OpenStudy (zzr0ck3r):

do you see why?

OpenStudy (zzr0ck3r):

f(x+2) shifts left and f(x-2) shifts right, horizontal shifts are counter intuitive...

OpenStudy (anonymous):

right, translations through the parent function

OpenStudy (zzr0ck3r):

yes, your parent function is 1/x

OpenStudy (zzr0ck3r):

ok so how far has it moved up?

OpenStudy (anonymous):

stretch of 5, or is that a reflection

OpenStudy (zzr0ck3r):

its a vertical shift of 5

OpenStudy (zzr0ck3r):

so we have (a/x+2) +5

OpenStudy (zzr0ck3r):

notice that horizontal shifts are counter intuitive but vertical shifts are exactly how one would expect.

OpenStudy (anonymous):

a=7?? Are u sure the equation doesn't having something to do w/ horizontal and vertical asymptotes instead of parental functions, see this is what i mean about me confusion to writing this equation

OpenStudy (zzr0ck3r):

a = 4

OpenStudy (zzr0ck3r):

7 = a/2+5 about your other question. we can think of it that way as well we see that the horizontal asymptote is at y = 5

OpenStudy (anonymous):

not to butt in, but since there is a "vertical asymptote" at x = -3, shouldn't the denominator be x + 3 ?

OpenStudy (zzr0ck3r):

its at -2

OpenStudy (anonymous):

oops sorry

OpenStudy (zzr0ck3r):

ive told you how to write the equation, you will get the same answer the other way. if you would like me to explain I will....

OpenStudy (anonymous):

or is it parental function i am confused... @zzr0ck3r

OpenStudy (anonymous):

yes please explain :)

OpenStudy (zzr0ck3r):

so you have the equation a/(x-h)+k if we let x get big we see that you will be very close to y=5 notice what happens in a/(x-h)+k if we let x get big we get a/(huge number)+k if we let x go to infinity we get a/infinity + k, so we get 0+k so for you k=5 now as far as h, we cant divide by 0 and this is what causes a vertical asymptote, so you look at your asymptote and see it occurs as x=-2 so if we let x=-2 then we have a/(x+2)+5, note that if you let x = 2 then you would be dividing by 0 and thus getting a vertical asymptote. to obtain a, you do the same thing as last time, it has nothing to do with asymptotes. so at first I showed you how to get the answer from translations, then I showed you how to get the answer from investigating asymptotes. both the same answer

OpenStudy (anonymous):

Got it so the equation can be found both ways, so let me see if i got it Would this be equation? y=4/(x+2)+5

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

Thanks so much!!!!!

OpenStudy (anonymous):

@satellite73 is this right can u double-check! :)

OpenStudy (zzr0ck3r):

if you understand it you should know its right, if you don't you should ask me a question

OpenStudy (anonymous):

No i think its right based on what i know but i struggle w/ math lol

OpenStudy (zzr0ck3r):

im not worried about being right, im worried you dont understand it